OpenStudy (anonymous):

on dividing a certain number by 342, we get 47 as remainder. if the same number is divided by 18, what will be the remainder?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

How you do this is first take any number that satisfies the first condition. So you need a number that when you divide it by 342 you have 47 left over. Since 389 = 342 + 47, we can use this number. Now, divide 389 by 18. We get 389/18 = 21 r 11. This remainder will be the same regardless of the original number we picked. To show this fact, try 731 = 2(342) + 47. 731/18 = 40 r 11...